if cos(theta)=7/25 , find the remaining two trig functions of theta and the measure of angle theta.
Answer by amarjeeth123(570)

cos(theta)=7/25
We have sin(theta)=√1-(cos(theta))^2=√1-(7/25)^2=√1-49/625=√576/625=24/25
tan(theta)=sin(theta)/cos(theta)=(24/25)/(7/25)=24/7
Solving we get theta=73.74 degrees.